OpenStudy (anonymous):

Points A,B, and C are collinear. Point B is between A and C. Solve for x AC=3x+3, AB= -1+2, and BC= 11

OpenStudy (jdoe0001):

thanks for sharing that

OpenStudy (jdoe0001):

AB = -1+2?

OpenStudy (anonymous):

-1+2x

OpenStudy (jdoe0001):

anyhow  \(\begin{cases} AC=3x+3\\ AB=-1+2x\\ BC=11 \end{cases} \\ \quad \\ AC-AB=BC\implies (3x+3)-(-1+2x)=11\) solve for "x"

OpenStudy (jdoe0001):



OpenStudy (mathmate):

Are you sure you didn't miss anything in AB=-1+2 ? It doesn't look right to me. It would have been written AB=1 @Ryan00